OpenStudy (anonymous):

So if I have the problem 1/4y+5=1/7y do I multiply everything by 7?

OpenStudy (anonymous):

you could however that wouldn't get you nice coefficients

OpenStudy (anonymous):

go ahead though

OpenStudy (anonymous):

actually do it because it will make the substraction easier

OpenStudy (anonymous):

I don't get how to do these ! I keep getting them wrong :,(

OpenStudy (anonymous):

so you'll get 7/4y +35 = y

OpenStudy (anonymous):

did you get that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now you need to get your y terms on one side

OpenStudy (anonymous):

then I subtract y from both sides?

OpenStudy (anonymous):

you can either minus the 1y and minus the 35 or you can just minus 7/4y

OpenStudy (anonymous):

basically you want to load one side with y and the other just be a constant or a number

OpenStudy (anonymous):

i would minus 7/4 y because then you wont have to do two steps

OpenStudy (anonymous):

so you should get 35=1y-7/4y

OpenStudy (ybarrap):

Multiply both sides by y then you get the simpler: 1/4 + 5y = 1/7 to get 5y = 1/7 - 1/4 = (4-7)/28 = -3/28 Ans. y = -3/(5*28)

OpenStudy (anonymous):

-3/4y= 35

OpenStudy (anonymous):

35*-4/3 = y

OpenStudy (anonymous):

So the answers 35-4/3=y ?

OpenStudy (anonymous):

*

OpenStudy (anonymous):

are the y's in the denominator or in the numerator

OpenStudy (anonymous):

there not in either there to the side of the fractions

OpenStudy (anonymous):

so its 1/4 y +5= 1/7 y

OpenStudy (anonymous):

so it is \[\frac{1}{4}y+5=\frac{1}{7}y\]

OpenStudy (anonymous):

Yes my love !

OpenStudy (anonymous):

personally i like using whole numbers instead of fractions so i would use the least common multiple of 4 and 7 which is 28 so i'd multiply everything by 28 and get: 7y + 140 = 4 y and then subtract 4y from both sides: 7y-4y + 140 = 4y - 4y 3y + 140 = 0 then subtract 140 from both sides 3y+140-140=-140 3y=-140 divide both sides by 3 3y/3=-140/3 y=-140/3

OpenStudy (anonymous):

your answer will be -140/3

OpenStudy (anonymous):

which is simply